# CulturalBench: a Robust, Diverse and Challenging Benchmark on Measuring the

### v1

@@ 2 Related Work
Benchmark # Annotators per Qn ($\uparrow$) Verified Qn Coverage (Verified #/Total #) ($\uparrow$) Data Filtering by Majority Votes Topic Inclusion # Topic ($\uparrow$) Source Best Model Performance ($\downarrow$)
Candle (Nguyen et al., 2022) 3 $0\%$ (0/1.1M) Predefined set 6 Web $81.4\%$ (GPT-3)
CultureAltas (Fung et al., 2024) 5 $0\%$ (0/10K) Predefined set 8 Wiki + LLM $93.1\%$ (GPT-3.5)
Normad (Rao et al., 2024) 2 $18.5\%$ (480/2.6K) Predefined set 4 Web + LLM $87.6\%$ (GPT-4)
@@
Blend (Myung et al., 2024) 5 $0\%$ (0/500) Discovery- based 5 Human + LLM $85.5\%$ (GPT-4)
@@
CulturalBench (Our Work) 5 $100\%$ (1227/1227) Discovery- based 17 Human + LLM $61.5\%$ (GPT-4o) (Human: $92.6\%$)
@@
Table 1: Comparison of existing cultural benchmarks on three criteria. Relative to existing benchmarks, CulturalBench is robust, diverse and challenging. Verified Qn Coverage refers to the human quality checks on the final collected questions on the benchmark, rather than intermediate steps of data collection. Best Model Performance refers to the average accuracy/F1 scores attained by best performing model on benchmark, with the model in parenthesis.
@@
Multicultural knowledge evaluation of LLMs have been widely investigated through building extensive knowledge bases (Shi et al., 2024; Keleg & Magdy, 2023); using socio-cultural surveys like World Value Survey (Durmus et al., 2023a; Tao et al., 2023; Ramezani & Xu, 2023); and generating more training data (Li et al., 2024). Here, we select four representative benchmarks with comparable model evaluation results, highlighting their limitations and the gaps that our CulturalBench aims to fill in Table 1.
Insufficient Quality Verification
@@
Existing cultural benchmarks usually conduct quality check during the intermediate steps on data collection such as the relevance of web-scraped knowledge (Fung et al., 2024), commonality of knowledge (Nguyen et al., 2022). Blend asked humans to directly curate answers and aggregating those inputs to form questions but did not verify the final questions by humans (Myung et al., 2024). Normad verified part of the rule-of-thumbs but with two humans only (Rao et al., 2024). As cultural knowledge is not easily verifiable for correctness, it is essential to have reliable annotations on final set of questions (as given to LLMs) by having expert human verification on the full set of questions and then filtering out questions that does not reach consensus.
Poor diversity of topics
@@
Many benchmarks have topics predefined prior to data collection, meaning that they are unlikely to fully capture the multi-faceted natured of cultural knowledge (Adilazuarda et al., 2024). Many prior works topics focus on narrow topics such as food (Nguyen et al., 2022), dating (Fung et al., 2024), social etiquette like dining (Palta & Rudinger, 2023; Dwivedi et al., 2023), visiting (Rao et al., 2024), and special elements in wider society like religions (Nguyen et al., 2022). To the best of our knowledge, only Blend uses a discovery-based approach to ask annotators to include all topics they believe to be relevant to culture(Myung et al., 2024), without restricting it to particular topics. CulturalBench extends this discovery-based approach helping us to identify diverse topics outlined in Fig. 3.
Over-reliance on Web Sources
@@
Existing benchmarks often rely on web sources directly such as web corpus (Nguyen et al., 2022), Wikipedia (Naous et al., 2023), and incorporated with LLMs’ generation (Rao et al., 2024; Fung et al., 2024). These non-human written benchmarks may not be challenging since the scraped web sources may be used during models pretraining (Petroni et al., 2019) and LLM generations may inherit the potential cultural bias (Arora et al., 2022; Cao et al., 2023; Liu et al., 2024). Given the performances of best-performing models ranging from 81.4% to 93.1% in the existing benchmarks in Table 1, those benchmarks are likely not sufficiently challenging for modern frontier LLMs. Our proposed CulturalBench is substantially more difficult with the best model (GPT-4o) only reaching 61.5% despite humans reaching 92.6%.
3 Data collection pipeline
Our data collection pipeline consists of three steps, as illustrated in Fig. 2: (1) Data collection via AI-assisted red-teaming (2) Human quality check on full data (3) Filtering with majority vote. Such a multi-step process enables us to collect robust data for CulturalBench.
3.1 Step 1: Data collection via interactive AI-assisted red teaming
Figure 2: Step-by-Step details on Data Collection and Validation.
Question Formulation. Human annotators are instructed to brainstorm culturally relevant scenarios based on their personal experiences of their cultures (e.g., Singaporeans use tissue packet to reserve seats). A step-by-step guideline with detailed examples is provided to inspire them, as shown in Appendix H.
The AI helper bot then transforms the scenario into a structured question with four options, which the annotators can review and edit afterward.
Question Verification & Revision. Human annotators can use the formulated question as basis to challenge the AI verifier in our interactive platform. The platform provide further assistance in revising the questions to make it more challenging by offering various revision strategies along with drafted examples (e.g., “Negate the Question”), as shown in Appendix H.
Internal Filtering. After collecting over 2,600 questions, the researchers carefully reviewed and removed those that are not relevant to any countries/regions (e.g., Bangladesh, Peru), resulting in a filtered set of 2,486 cultural questions.
3.2 Step 2: Human Quality Check
Recruitment Criteria. We collected questions at the country/regional level, pairing each question with a specific region. To ensure culturally attuned and thorough verification, we recruited five annotators for each region through the Prolific platform  [^1] https://www.prolific.com. We set two main criteria to ensure that the recruited annotators have a deep understanding of the culture of the targeted country or region – (1) Nationality (2) Primary residence before age 18. For certain cultures (e.g. the United States, the United Kingdom), when the platform allowed more detailed selections and the collected question targeted specific groups in the country/region, we added detailed criteria such as ethnicity (e.g. African American, Native American), and place of residence (e.g., Wales).
Multiple Selection Settings. To better reflect the true representation of each cultural question, we allow annotators to select multiple answers on our questions with four options. As a result, some questions may have more than one majority-vote answer. This approach also helps test models’ mode-seeking behavior, examining whether they rely solely on cultural stereotypes (i.e., modes) without considering broader cultural diversity.
3.3 Step 3: Filtering by Majority Vote & Constructing Benchmarks
Majority Vote Criteria. To build a robust benchmark that captures the accurate representation on cultural knowledge, we set the majority-vote threshold to be $>=4$ out of 5 annotators. During human validation, we first filtered out questions without majority consensus, resulting in a final set of 1,227 questions. Subsequently, we further processed the remaining questions. To construct our CulturalBench in two setups (CulturalBench-Easy: Multiple-choice, CulturalBench-Hard: True/False), we processed the questions differently depending on the numbers of majority votes they contain.
(1) Single-Mode Questions (Only one majority vote). For CulturalBench-Easy, we directly keep the original question with four options. The gold label is the option with a majority vote (i.e., A, B, C or D). For CulturalBench-Hard, we transform the question with four options into four binary questions. For instance, the question drafted (e.g., “What do Singaporeans …? A. Tissue … D. …”) will form binary questions (e.g. “Is this answer true or false for this question? Answer True or False only. Question: What do Singaporeans …? Answer: Tissue.”).
(2) Multi-Mode Questions (More than one majority votes). For CulturalBench-Easy, we reframe the question to allow multiple statements. The four drafted options (e.g., “A. Tissue”) become the four statements in questions (e.g., “statement (i) Tissue”). To ensure the models know the possibility of questions containing multiple correct labels, we add the instruction on question directly with (“Select the options with all applicable statements”). For CulturalBench-Hard, we follow the same construction approach (transforming four options to four binary questions) as single-mode questions.
4 CulturalBench Description and Discovered Topics
4.1 Descriptive Statistics on CulturalBench
Our benchmarks cover a wide range of global regions, spanning 45 countries and regions, including underrepresented regions such as Bangladesh, Zimbabwe, and Peru. A detailed breakdown of regional distribution can be found in Appendix B while example questions by topic are available in Appendix C.
CulturalBench-Easy. It contains 1,227 multi-choice questions, each with four options. For instance, a question of “What do Singaporeans usually use to reserve seats?” with options of “A. Tissue … D. Book” as shown in Fig. 1. The gold label is the correct option (A, B, C or D). For multi-mode questions (i.e., questions with more than one answer), we added an instruction of “Selecting the option with all applicable statements” to ensure that models consider all possible answers for fair evaluation. For instance, a question of “What do Singaporeans…? Selecting the option … statements. i) Tissue … iv) Books” with options of “A. (i) … D. (i), (ii), (iv)”. The questions contain 17.2 words on average ($\sigma=$ 12.06). Options at various positions have similar number of whitespace-separated words on average, specifically option A with 5.48 words ($\sigma=$ 4.24), option B with 5.44 words ($\sigma=$ 4.27), option C with 5.57 words ($\sigma=$ 4.24), and option D with 5.57 words ($\sigma=$ 4.24).
CulturalBench-Hard. In this dataset, each question is transformed into four binary true/false questions, requiring models to evaluate each option separately. For example, the earlier multiple-choice example in CulturalBench-Easy will transform into four binary questions such as “Is this answer true or false for this question? Question: What do Singaporeans usually use to reserve seats? Answer: Tissue.”, as shown in Fig. 2. The gold label in this case is either True or False. This set contains 1,227 ×4$\times$ 4=× 4 = 4,912 True/False judgement questions. The questions contain 14.3 words ($\sigma=$ 5.27) and the answers contain 5.72 words ($\sigma=$ 4.21).
4.2 Diverse Topics Discovered Across Cultures
Most existing cultural benchmarks have predefined topics to collect data on, typically on universal topics such as dining (Adilazuarda et al., 2024). However, this approach can overlook cultural elements unique to specific regions. To capture a broader spectrum of cultural topics, we adopted a discovery-based approach by encouraging human annotators to brainstorm cultural concepts from their personal experiences. The detailed instruction for annotators can be found in Appendix H. CulturalBench spans a diverse range of cultural elements with 17 topics under three categories (Daily life, Social Etiquette, and Wider Society), as shown in Fig. 3. Daily life relates to the everyday experiences of people e.g., Workplace. Social Etiquette means the acceptable norms in society e.g., Greeting. Wider Society included special elements for broader spectrum of cultural topics e.g., Celebrations. We classified questions into topics by prompting GPT-4o-mini. The classification prompt and the topic detailed definitions are in Appendix C.
To collect diverse data for each culture, we allow each annotator to create at most 3-7 questions, depending on the availability of annotators for each region. Notably, in curating CulturalBench, we observed that people from different regions focused on distinct topics. For instance, annotators from Italy and Mexico provided more questions related to Food, with 15 out of 35 questions and 13 out of 49 questions respectively. In contrast, participants from South Africa and India focused more on Religion, contributing 19 out of 58 questions and 14 out of 46 questions respectively. Our discovery-based approach allow us to capture diverse cultural elements from people in different regions without being limited by a predefined set of topics.
5 Experiments: Evaluation of LLMs on CulturalBench
We evaluate 30 current LLMs in a zero-shot setting on CulturalBench in two setups: (1) CulturalBench-Easy: Multiple choice; (2) CulturalBench-Hard: True/False. We prompted the models to ensure they follow the output format to allow fair comparison. The detailed prompt is in Appendix D. To avoid exposing the correct answers to models for fair comparison, our annotation platform, which involves using OpenAI APIs did not allow the collected data to be used for further training.
CulturalBench-Easy. We evaluate model performance by measuring accuracy, specifically whether the model correctly identifies the label for each multiple-choice question. A random baseline can achieve 25%.
CulturalBench-Hard. We evaluate model performance based on the proportion of tasks in which the model can get all four options predicted correctly. For each task, an LLM has to make four binary judgements (True/False) from the transformation of four options in each multiple choice question. To demonstrate robust cultural knowledge, we believe the LLM has to accurately which option(s) are False as well as which option(s) are True. A random baseline can achieve $0.5^{4}$ = 6.25%.
5.1 Comparing LLMs on two benchmarks across model family and size
Figure 4: Models performance on CulturalBench-Hard with random baseline at 6.25% and human performance at 92.6%.
We show the performance of 18 models across model families and sizes on CulturalBench-Hard in Fig. 4. The corresponding Fig for CulturalBench-Easy is in Appendix A.
Models performance on CulturalBench-Easy. The best-performing model, GPT-4o, achieves 88.8% accuracy, slightly lagging behind human performance at 92.4%, as illustrated in Appendix A. Nonetheless, this benchmark remains an effective tool for assessing model capabilities, with the lowest score of 61.7% (Claude 3-Haiku) clearly highlighting the wide range of model performance.
Models performance on CulturalBench-Hard. As shown in Fig. 4, this benchmark is significantly more challenging for current LLMs, with accuracy ranging from 21.4% for Llama3-8b to 61.5% for GPT-4o. These scores are considerably lower compared to the human baseline of 92.6%, highlighting the difficulty of the task even for the most advanced models.
Models performance improves as model size increases. In Fig. 4, we present the performance of models from six different families, such as GPT, Llama, and Qwen. Overall, the results demonstrate a trend of improved performance as model size increases. For example, within the Claude-3 family, the models show a clear progression in accuracy: Claude 3-Haiku achieves 25.3%, Claude 3-Sonnet reaches 34.5%, and Claude 3-Opus attains 52.9%. This pattern is consistent across most of the model families, indicating that larger models generally perform better on our CulturalBench-Hard.
Why do the two setups on CulturalBench have such model performance difference? We hypothesize that the models can guess for the most possible answer on CulturalBench-Easy under the multiple-choice setting. We compute the embedding for the country name and separately for each option using OpenAI text-embedding-3-small. By using a simple heuristic of choosing the option with highest cosine similarity with the country name (e.g. Bangladesh), we attain 40.42% accuracy. This is intriguing as it is substantially above the random baseline (25%), without needing considering the question at all. We find that the cosine similarity difference between the correct option and the country name is significantly higher than the difference between options average and the country (0.166 vs. 0.145; Kruskal-Wallis $p$$\leq$0.01). This shows the possibility of models guessing based on one (out of many possible) heuristics in multiple-choice setup without understanding (or even knowing) the question. This stresses the importance on using the binary (True/False) for each of the four options per question in CulturalBench-Hard to accurately assess cultural knowledge of LLMs.
5.2 Investigating effects of question type and time version of models
Figure 5: Analysis on question type (Left) and time version (Right). For question types, we demonstrate models struggle at answering questions with multi-modes (more than one correct answers). For time version, we show the improved performance of models across time.
LLMs show distinct gaps between question types, unlike humans. We evaluate the performance based on question types – (1) Single-mode (N=1141) and (2) Multi-mode (N=86). The first type refers to the questions with only one correct, majority-voted answer while the second type includes questions with multiple correct answers, as explained in Section 3.3. In Fig. 5 (Left), the average across all models shown in Fig. 4 is 43.4% on Single-mode questions and 25.2% on Multi-mode questions, revealing a significant gap of 18.2% between the two. Similarly, the best model (GPT-4o) exhibits a 19.8% performance difference between these question types. In contrast, human baselines show only a 2.2% difference, indicating that humans handle cultural diversity more effectively than models. This discrepancy suggests that models struggle to account for cultural nuances due to their mode-seeking tendencies, as discussed by (Tajwar et al., 2024).
Models in the same series improve across time versions. In Fig. 5 (Right), we evaluate four models (GPT-3.5-turbo, Claude Sonnet, GPT-4-turbo, and GPT-4o) across different available time versions (e.g., we evaluate GPT-3.5-turbo on ‘GPT-3.5-turbo-1106’ as earlier version and ‘GPT-3.5-turbo-0125’ as later version). Overall, all four models demonstrated an increasing trend in performance across time. The largest improvement is shown in the Claude Sonnet model improves in performance from 34.5% (Claude-3 Sonnet) to 51.5% (Claude-3.5 Sonnet). By comparison, strongest model (GPT-4o) shows only a modest 0.5% increase between versions.
5.3 Studying different providers’ LLMs on questions from different regions
Figure 6: Models performance (18 models tested from Fig. 4) by different providers on different cultural groups. We further compare five representative models (GPT-4o, Claude Sonnet, Mistral Large, Llama-3.1 405b and Qwen-2 72b) from different model families.
We include detailed performance of models across different family and sizes (shown in Fig 4) to understand how well different models performance in questions relating to different geographic regions at a continent/sub-continent level.
Overall, models perform better in questions relating to North America, South Asia, and West/South Europe. From Fig. 4, it is evident that models achieve higher performance averages in regions like North America (58.0%), South Asia (52.3%), West Europe (47.1%) and South Europe (45.4%). We hypothesize that the higher performance in these regions can be attributed to several factors including their representation on web-data used for model training (Longpre et al., 2023) and the proportion of annotators recruited from these regions by LLM providers to curate post-training alignment data. For instance, many annotators are known to be recruited from India as they have good English ability and costs substantially less than their counterpart in the US (Lohchab & Roy, 2024).
Models score lower in questions relating to South America, East Europe, and the Middle East. Models exhibit lower performances on average in regions like South America (38.2%), East Europe (37.6%), and Middle East/West Asia (33.6%), compared to neighbouring regions such as North America (58.0%) and West Europe (47.1%). These disparities suggest insufficient representation of cultural knowledge from these regions in the training data.
GPT-4o leads in most regions, followed by Llama-3.1 405b and Claude-3.5 Sonnet. GPT-4o consistently ranks highest across most regions among all tested models. Llama-3.1 405b shows strength in regions where cultural knowledge is traditionally less represented, such as South America and East Europe, while Claude-3.5 Sonnet performs particularly well in other regions e.g., Oceania, West Europe, Africa, and Southeast Asia.
Chinese Model Providers (Qwen-2-72b) and European Model Providers (Mistral Large) are not stronger in cultural knowledge relating in their region. Despite claims of specialization in local languages, Qwen-2-72b and Mistral Large do not outperform other models in their respective regions in terms of cultural knowledge. For example, Qwen-2-72b scores 50.7% on East Asia, while GPT-4o achieves 61.4%. Similarly, Mistral Large underperforms in West Europe (48.9%) compared to GPT-4o (54.3%). These results suggest that local language proficiency alone is not sufficient for strong cultural competence.
6 Limitations and future work
@@
Multilingual vs. Multicultural. We develop an English-only benchmark as the initial step in evaluating models’ cultural knowledge. This approach facilitates fair comparisons of cultural understanding across different regions. For instance, in underrepresented regions such as Bangladesh, the availability of training data in local languages is often limited. As a result, models lacking sufficient exposure to these languages may struggle to comprehend questions phrased in them (Yong et al., 2023). By employing an English-only benchmark, we can assess models’ cultural knowledge regarding these underrepresented areas without considering their (lack of) proficiency in low-resource languages. Additionally, prior research on multilingual models’ emotional understanding (Havaldar et al., 2023) and reasoning skills (Liu et al., 2023) indicates that a model’s multilingual capabilities may not necessarily correlate with its multicultural competencies. Notably, our discovery-based benchmark includes language elements on some questions, particularly in the Language and Communication topic with 6.2%. For example, we included questions like: “What do Singaporeans usually say at the end? A. lah …”. As we await advances in developing stronger multilingual abilities in models for low-resource languages, our goal is to establish a robust, diverse, and challenging benchmark to track our progress toward addressing the uneven representation of cultural knowledge.
@@
Small sample of human verifiers on subjective cultural knowledge. Due to the limitations of crowd-sourcing platforms like Prolific, the number of available annotators from underrepresented regions, such as Bangladesh, is quite small (fewer than 30 active human annotators). As a result, we were able to recruit only five annotators for consistency verification. To enhance the robustness of our dataset, we allow human verifiers to select multiple labels for each question, ensuring that all possible answers are captured. Additionally, we establish a strict majority-vote threshold (majority votes $\geq$ 4 out of 5). During the annotation process, we also provide two extra options: “I don’t have knowledge” and “This question is unanswerable” – to enable annotators to indicate when they cannot provide a response.
@@
Further fine-grained culture classification. We noticed that the country/region classification adopted by our CulturalBench may not capture the cultural diversity within each region. However, the data annotation platform we accessed does not have a further fine-grain classification when recruiting human annotators for most of the regions except for the United States and the United Kingdom. To capture the diversity on these two countries, we revisited the data that have been filtered by having not enough majority votes and with mostly responses of “I don’t have knowledge”. For example, questions asking for the Welsh custom in the United Kingdom may not be answerable for people living in England. Then, we conducted a second round of human quality check by assigning those questions for the specific groups of human annotators (e.g., people living in Wales in the United Kingdom), as explained in Section 3.2. We hope to see more data annotation tools for different local cultures to facilitate more fine-grained cultural data collection.
Strong instruction prompts and strict evaluation criteria on models’ outputs. We evaluated models on zero-shot setting with the prompts. However, for some models such as Claude-3 Haiku, they need more instructions to have the right formatting. Therefore, we have added one-line instruction for all models to ensure they outputting the answer in the correct format on our evaluation, as described in Appendix D. However, with the strong instruction prompt, sometimes they still refuse to answer the questions e.g., “This question …” rather than outputting the four options (i.e., A, B, C, or D) on CulturalBench-Easy or the binary labels (i.e., True/False) on CulturalBench-Hard. To ensure the fair evaluation, we set the output token to be 2. The model is treated as answering correctly when its output contains the correct labels only.
7 Conclusion
We present CulturalBench in two setups: CulturalBench-Easy and CulturalBench-Hard. By establishing a robust, diverse, and challenging benchmark to track our progress in cultural knowledge, we hope it can motivate LLM providers to develop models that can be helpful to users across more geographical regions.
Ethics statement
@@ 7 Conclusion
Our data collection has been reviewed by university’s IRB board to ensure it has no harm on human annotators. We pay annotators according to our vendor (Prolific)’s guidance, which is higher than the local wage requirement. Our annotation guidance has specifically asked annotators to not include their personal identifiable information when giving their responses. Before human verification, our internal team has reviewed the collected data to ensure there is no harmful or unsafe context such as sexual or violence content.

### v2

@@ 5 Related Work
Benchmark Task # Annotators per Qn ($\uparrow$) Verified Qn Coverage (Verified #/Total #) ($\uparrow$) Data Filtering by Majority Votes Topic Inclusion # Topic ($\uparrow$) Source Best Model Performance ($\downarrow$)
FORK (Palta and Rudinger, 2023a) Text-based short-form QA 2 100% (184/184) Predefined set 1 Human 74.7% (Average for BERT series models)
BERTAQA (Etxaniz et al., 2024) Text-based short-form QA NA NA NA NA 12 Web 91.7% (GPT-4)
CVQA (Romero et al., 2024) Visual QA 2 (including author of Q) 100% (10,374/10,374) Predefined set 10 Humans 75.4% (GPT-4o)
NormAd (Rao et al., 2024) Text-based short-form QA 2 18.5% (480/2.6K) Predefined set 4 Web + LLM 87.6% (GPT-4)
@@
Blend (Myung et al., 2024) Text-based short- form/long-form QA 5 0% (0/500) Discovery- based 5 Human + LLM 85.5% (GPT-4)
@@
CULTURALBENCH (Our Work) Text-based short-form QA 5 100% (1696/1696) Discovery- based 17 Human + LLM 61.4% (o1)
@@
Table 1: Comparison of existing cultural benchmarks on three criteria. Relative to existing benchmarks, CulturalBench is robust, diverse and challenging. Verified Qn Coverage refers to the human quality checks on the final collected questions on the benchmark, rather than intermediate steps of data collection. Best Model Performance refers to the average scores attained by best performing model on benchmark, with the model in parenthesis. Reported metrics are as follows: Candle: Precision; CultureAltas: F1; Normad: accuracy; Blend: accuracy; CulturalBench: accuracy.
@@
Multicultural knowledge evaluation of LMs has been widely investigated through building extensive knowledge bases (Shi et al., 2024; Keleg and Magdy, 2023); using socio-cultural surveys like World Value Survey (Durmus et al., 2023a; Tao et al., 2023; Ramezani and Xu, 2023); and generating more training data (Li et al., 2024). Here, we select four representative benchmarks with comparable model evaluation results, highlighting their limitations and the gaps that our CulturalBench aims to fill in Table 1.
Insufficient Quality Verification.
@@
Existing cultural benchmarks usually conduct quality checks during the intermediate steps on data collection such as the relevance of web-scraped knowledge (Fung et al., 2024), commonality of knowledge (Nguyen et al., 2022). Blend asked humans to directly curate answers and aggregating those inputs to form questions but did not verify the final questions by humans (Myung et al., 2024). Normad verified part of the rule-of-thumbs but with two humans only (Rao et al., 2024). As cultural knowledge is not easily verifiable for correctness, it is essential to have reliable annotations on the final set of questions (as given to LMs) by having expert human verification on the full set of questions and filtering out questions without consensus answers.
Predefined cultural topics.
@@
Many benchmarks have topics predefined prior to data collection, meaning that they are unlikely to fully capture the multi-faceted natured of cultural knowledge (Adilazuarda et al., 2024). Many prior works topics focus on narrow topics such as food (Nguyen et al., 2022), dating (Fung et al., 2024), social etiquette like dining (Palta and Rudinger, 2023b; Dwivedi et al., 2023), visiting (Rao et al., 2024), and special elements in wider society like religions (Nguyen et al., 2022). While Blend uses a discovery-based approach (Myung et al., 2024), CulturalBench extends it to identify more diverse topics (6 vs. 17) with details in Fig. 6.
Over-reliance on Web Sources.
@@
Existing benchmarks often rely on web sources directly such as web corpus (Nguyen et al., 2022), Wikipedia (Naous et al., 2023), and incorporated with LMs’ generation (Rao et al., 2024; Fung et al., 2024). These non-human written benchmarks may not be challenging since the scraped web sources may be used during models pretraining (Petroni et al., 2019) and LM generations may inherit potential cultural biases (Arora et al., 2022; Cao et al., 2023; Liu et al., 2024). Given the highest performance of models range from 81.4% to 93.1% in the existing benchmarks by relatively weaker models (e.g. GPT-3/3.5/4) in Table 1, those benchmarks are likely not sufficiently challenging for modern frontier LMs (e.g. o1). Our proposed CulturalBench is substantially more difficult with the best model (OpenAI o1) only reaching 61.5%, far from the human performance of 92.4%
6 Conclusion
Inspired by the human-AI red teaming, we establish CulturalBench: a robust, diverse, and challenging benchmark for measuring LMs’ culture knowledge. We hope the community can build upon our work to accelerate our progress in improving cultural knowledge of LMs.
Limitations
@@ 6 Conclusion
We develop an English-only benchmark as the initial step in evaluating models’ cultural knowledge. This approach facilitates fair comparisons of cultural understanding across different regions. For instance, in underrepresented regions such as Bangladesh, the availability of training data in local languages is often limited. As a result, models lacking sufficient exposure to these languages may struggle to comprehend questions phrased in them (Yong et al., 2023). By employing an English-only benchmark, we can assess models’ cultural knowledge regarding these underrepresented areas without considering their (lack of) proficiency in low-resource languages. Additionally, prior research on multilingual models’ emotional understanding (Havaldar et al., 2023) and reasoning skills (Liu et al., 2023) indicates that a model’s multilingual capabilities may not necessarily correlate with its multicultural competencies.
Small sample of human verifiers on subjective cultural knowledge.
@@
Due to the limitations of crowd-sourcing platforms like Prolific, the number of available annotators from underrepresented regions, such as Bangladesh, is quite small (fewer than 30 active human annotators). As a result, we were able to recruit only five annotators for consistency verification. To enhance the robustness of our dataset, we allow human verifiers to select multiple labels for each question, ensuring that all possible answers are captured. Additionally, we establish a strict majority-vote threshold (majority votes $\geq$ 4 out of 5). During the annotation process, we also provide two extra options: “I don’t have knowledge” and “This question is unanswerable” – to enable annotators to indicate when they cannot provide a response, as illustrated in Appendix Fig. 13 and 14.
Further fine-grained culture classification.
@@
We noticed that the country/region classification adopted by our CulturalBench may not capture the cultural diversity within each region. However, the data annotation platform we accessed does not have a further fine-grain classification when recruiting human annotators for most of the regions except for the United States and the United Kingdom. To capture the diversity on these two countries, we revisited the data that have been filtered by having not enough majority votes and with mostly responses of “I don’t have knowledge”. For example, questions asking for the Welsh custom in the United Kingdom may not be answerable for people living in England. Then, we conducted a second round of human quality check by assigning those questions for the specific groups of human annotators (e.g., people living in Wales in the United Kingdom), as explained in Section 2.2. We hope to see more data annotation tools for different local cultures to facilitate more fine-grained cultural data collection.
Risk
Our benchmark may unintentionally increase existing biases in LM Cultural Knowledge evaluation. Even through we tried our best to capture cultural questions from different regions, some cultural perspectives might be overrepresented or underrepresented (due to the availability/perspectives of annotators on the Prolific platform), leading to skewed evaluations.
Ethical Considerations
@@
Our data collection has been reviewed by the university’s IRB board to ensure it has no harm on human annotators. We pay annotators according to our vendor (Prolific)’s guidance, which is higher than the local wage requirement. Before annotation, we explain the annotation task, and how their data will be collected and transformed into the quiz questions for models. With such understanding, they consented to their annotations being used in such a manner. Our annotation guidance has specifically asked annotators to not include their personal identifiable information when giving their responses. Before human verification, our internal team has reviewed the collected data to ensure there is no harmful or unsafe context such as sexual or violence content.